if a and b do a job together in two hours. Working alone A does the job in 5 hours. how long will it take B to do the job alone?
Answer by richwmiller(17219)

rate for A alone 5
rate for A & B together 2
1/a+1/b=1/x
1/5+1/b=1/2
1/b=1/2-1/5
1/b=5/10-2/10
1/b=3/10
1/b=1/3.33333333
b=3.33333333
rate for B alone 3.33333333
check
1/5+1/3.33333333=1/2
3.33333333/16.6666667+5/16.6666667=1/2
8.33333333/16.6666667=1/2
ok
codewa